# UMA: Ultra-detailed Human Avatars via Multi-level Surface Alignment

### v1

@@ 3.1 Gaussian-based Avatar Representation
Animatable Gaussian Textures. Building upon the drivable template mesh, we model the fine-grained and motion-aware appearance and geometry of the clothed human using Gaussian textures $\mathbf{T}^{\mathrm{gs}}_{f}\in\mathbb{R}^{N\times 62}$  [7] in the template’s UV space. Each of the $N$ texels covered by a triangle stores the parameters of a 3D Gaussian splat $(\bm{\bar{\mu}}^{\mathrm{uv}}_{i},\mathbf{\bar{d}}^{\mathrm{uv}}_{i},\mathbf{q}^{\mathrm{uv}}_{i},\mathbf{s}^{\mathrm{uv}}_{i},\mathbf{\alpha}^{\mathrm{uv}}_{i},\bm{\eta}^{\mathrm{uv}}_{i})_{f}\in\mathbb{R}^{62}$ where $i$ denotes the $i$th texel. Notably, the canonical Gaussian position $\bm{\bar{\mu}}_{\mathrm{uv},i}$ is derived from the non-rigidly deformed template mesh $\bar{\mathbf{V}}_{f}$ through barycentric interpolation. To model the finer-level details, i.e., wrinkles, of the dynamic clothed human, a learnable motion-aware offset $\mathbf{\bar{d}}_{\mathrm{uv},i}$ is applied to each Gaussian Splat in the canonical space. Similar to the drivable template mesh, the Gaussian splats can be pose from the deformed canonical position $(\bm{\bar{\mu}}_{i}+\mathbf{\bar{d}}_{\mathrm{uv},i})$ to the posed space $\bm{\mu}_{i}$, through Dual Quaternion skinning  [62]. The remaining parameters denote the rotation quaternion $\mathbf{q}^{\mathrm{uv}}_{i}$, anisotropic scaling $\mathbf{s}^{\mathrm{uv}}_{i}$, opacity $\mathbf{\alpha}^{\mathrm{uv}}_{i}$, and spherical harmonics coefficients $\bm{\eta}^{\mathrm{uv}}_{i}$.
@@
Predicting Motion-aware Gaussians. Thanks to the texel-based parameterization, the mapping between the skeletal motion $\bm{\bar{\theta}}_{f}$ and motion-aware Gaussian splats $\mathbf{T}^{\mathrm{gs}}_{f}\in\mathbb{R}^{N\times 62}$ can be formulated as an image-to-image translation task [63]. Therefore, the surface is a learned function of skeletal pose, which is encoded as positional $\mathbf{T}^{\mathrm{p}}_{f}$ and normal textures $\mathbf{T}^{\mathrm{n}}_{f}$ of $\mathbf{V}_{f}$. Then, two separate convolutional decoders $\mathcal{E}_{\mathrm{app}}(\mathbf{T}^{\mathrm{n}}_{f},\mathbf{T}^{\mathrm{p}}_{f})$ and $\mathcal{E}_{\mathrm{geo}}(\mathbf{T}^{\mathrm{n}}_{f},\mathbf{T}^{\mathrm{p}}_{f})$ regress $(\mathbf{\bar{d}}^{\mathrm{uv}}_{i},\mathbf{q}^{\mathrm{uv}}_{i},\mathbf{s}^{\mathrm{uv}}_{i},\mathbf{\alpha}^{\mathrm{uv}}_{i})_{f}$ and $(\bm{\eta}^{\mathrm{uv}}_{i})_{f}$, respectively.
@@
Splatting Gaussians to Image Space. We adopt Analytical Splatting [64] $R^{\mathrm{analytical}}(\mathbf{T}^{\mathrm{gs}}_{f},\mathbf{C}_{c})=I^{\mathrm{R}}_{f,c}$ to render the final image $I^{\mathrm{R}}_{c}$ by integration over the actual pixel size during the splatting. This provides improved anti-aliasing over the original 3DGS tile-based rasterizer, which is particularly useful when supervising on ultra-resolution data, i.e., 6K resolution, and when focusing on detail preservation – the goal of this work.
@@
$\footnotesize\mathcal{L}_{\mathrm{gau}}(I^{\mathrm{R}}_{f,c})=\mathcal{L}_{1}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c})+\mathcal{L}_{\mathrm{ssim}}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c}),+\mathcal{L}_{\mathrm{mrf}}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c})$ (5)
consisting of image-based $\mathcal{L}_{1}$, structural $\mathcal{L}_{\mathrm{ssim}}$, and perceptual losses $\mathcal{L}_{\mathrm{mrf}}$  [65].
3.2 Key Technical Challenge
@@
Misalignment in Depth. In the context of animatable clothed human avatars, the highly diverse clothing dynamics cannot be fully explained by skeletal motions alone, as similar skeletal poses may result in significantly different surface deformations, often referred to as the one-to-many mapping issue [8]. As a result, when using Eq. 4 to supervise the avatar’s geometry (Eq. 1), the regression-based approach predicts an average surface that fits the multiple possible surface deformations in a least-squares sense due to this one-to-many mapping despite being supervised on accurate ground truth geometry. This manifests in misalignment in depth, which hinders accurate supervision of fine-grained appearance details, i.e., the motion-aware Gaussian Splats $\mathbf{T}^{\mathrm{gs}}_{f}$, since Gaussians may be wrongly projected onto training images (see Fig. 3 a)).
On-surface Drift. Even if depth misalignment is resolved, the ground-truth geometry obtained from implicit-based methods [61] still lacks temporal correspondence among frames and correspondence between the template surface and the ground truth surface. Thus, the second source of error is noticeable drift on the surface between the drivable template mesh and the ground-truth surface since the Chamfer loss can be minimized even if the correspondence is wrong. For example in Fig. 3 b), the blue Gaussian is never mapped onto the correct location (marked in blue). Nonetheless, the Chamfer loss is minimized. Even worse, the Gaussian is mapped onto different ground truth surface locations for different frames. Thus, the rendering losses will assign different colors for different frames to the same Gaussian leading to a blurred average.
@@ 3.3 Depth Alignment
Figure 3: Key Technical Challenges. (a) The depth misalignment leads to gradient conflicts when supervised with multi-view images, like the blue splat. (b) Even if the depth misalignment is resolved, the surface drift between different frames results in gradient conflicts, resulting in averaged and blurry appearance.
@@
To address the misalignment in depth caused by the one-to-many mapping between the skeletal motion and the surface geometry, we introduce a learnable per-frame latent
$\mathcal{H}_{\mathrm{lat}}(e(f/f_{\mathrm{max}}))=\mathbf{z}_{f}\in\mathbb{R}^{16}$ (6)
@@
as an additional input alongside the skeletal motion $\bm{\bar{\theta}}_{f}$ for predicting the drivable template mesh. This latent accounts for the stochasticity and effectively establishes a one-to-one mapping. Here, $f_{\mathrm{max}}$ denotes the maximum number of training frames, $e(\cdot)$ is a positional encoding [66], and $\mathcal{H}_{\mathrm{lat}}$ is a shallow MLP. We then update our deformable mesh (Eq. 1) by concatenating the learnable per-frame latent code $\bm{z}_{t}$ channel-wise to the input graph nodes of, both, the embedded deformation network $\mathcal{F}_{\mathrm{eg}}(\bar{\bm{\theta}}_{f},\mathbf{z}_{f})$ and the per-vertex deformation network $\mathcal{F}_{\mathrm{delta}}(\bar{\bm{\theta}}_{f},\mathbf{z}_{f})$.
@@
We refer the drivable template mesh as $M(\bm{\bar{\theta}}_{f},\mathbf{z}_{f})$ – now also taking the latent as parameter – and then train the drivable human avatar as before using the losses defined in Eq. 4. At inference, we set $\mathbf{z}_{f}$ to 00 to drive the character with novel motions. Please also refer to the supplemental materials for more details.
3.4 Vertex-level Alignment
@@
Through the depth alignment (Sec. 3.3), the gap between the posed template mesh $\mathbf{V}_{f}$ and the ground truth surface $\mathbf{V}^{\mathrm{gt}}_{f}$ is significantly reduced by addressing the motion ambiguities. However, since the ground truth reconstruction $\mathbf{V}^{\mathrm{gt}}_{f}$ inherently lacks correspondence over time, surface drift is still inevitable, which will result in blurred appearance and geometry, as discussed in Sec. 3.2. Some prior works [10] propose to additionally supervise the geometric deformation with a differentiable image rendering loss to account for the drifting. However, we found that such loss is easily stuck in local minima while not being robust to shading and change in lighting (see also Sec. 4.2). Instead, our solution builds up on recent advances on foundational 2D point tracking methods [11] that offer a promising alternative.
@@
2D Point Tracking. A 2D point tracker [11] estimates 2D point tracks between a pair of images $(\mathbf{I},\mathbf{I}^{\prime})$, e.g., consecutive frames of a video. It can be defined as the function
$c(\mathbf{I},\mathbf{I}^{\prime},\mathbf{p})=\mathbf{p}^{\prime}\in\mathbb{R}^{2}$ (7)
@@
taking the image pair as well as a 2D image coordinate $\mathbf{p}\in\mathbb{R}^{2}$ (of image $I$) and estimates the corresponding matching point $\mathbf{p}^{\prime}$ in image $I^{\prime}$. Since those trackers are trained on large-scale real world data through regression, they are typically robust to lighting changes and do not suffer from local optima. However, our goal is to carefully align our drivable template $M(\bm{\bar{\theta}}_{f},\mathbf{z}_{f})$ over a long multi-view sequence, which is not natively supported by those trackers.
@@
Render-to-image 2D Point Tracks. Our key idea is to compute 2D correspondence between the rendering $\mathbf{I}^{\mathrm{rast}}_{c,f}$ of the deformed template $\mathbf{V}_{f}$ observed from camera $c$ using a rasterizer [67] $R^{\mathrm{rast}}(\mathbf{V}_{f},\mathbf{T}_{\mathrm{0}},\mathbf{C}_{c})=\mathbf{I}^{\mathrm{rast}}_{f,c}$ and the training frame $\mathbf{I}_{c,f}$. Notably, as the deformed template suffers from drift, the rendering and the ground truth do not perfectly align and the point tracks $c(\mathbf{I}^{\mathrm{rast}}_{c,f},I_{f,c},\mathbf{p})$ will capture exactly this drift. To acquire the static texture $\mathbf{T}_{\mathrm{0}}$ of the deformable mesh, we perform texture unprojection [9] using the first frame of the multi-view video $\mathbf{I}_{c,0}$. Now, we can compute the drift for each vertex $\mathbf{V}_{f,i}\in\mathbb{R}^{3}$ in 2D image space as
$\small c(\mathbf{I}^{\mathrm{rast}}_{f,c},\mathbf{I}_{f,c},\mathbf{C}_{c}\mathbf{V}_{f,i})=c(\mathbf{I}^{\mathrm{rast}}_{f,c},\mathbf{I}_{f,c},\mathbf{v}_{f,c,i})=\mathbf{p}_{f,c,i}\in\mathbb{R}^{2}$ (8)
by querying $c(\cdot,\cdot,\cdot)$ at the vertex’ projected 2D position $\mathbf{v}_{c,f,i}\in\mathbb{R}^{2}$.
@@
3D Lifting and Consensus Finding. Next, we lift per-view 2D correspondences $\{\mathbf{p}_{f,c,i}\}_{c=1}^{C}$ of vertex $i$ to 3D correspondences $\{\mathbf{P}_{f,c,i}\}_{c=1}^{C}$ by querying depth maps defining $\mathbf{P}_{f,c,i}=\mathbf{C}^{-1}_{c}\mathbf{N}_{f,c}[\mathbf{p}_{f,c,i}]$ using bilinear interpolation []delimited-[][$\cdot$][ ⋅ ] and applying the inverse camera transform. To find consensus across views, we define a score function
$s_{f,c,i}=\frac{(\mathbf{V}_{f,i}-\mathbf{o}_{c})}{||\mathbf{V}_{f,i}-\mathbf{o}_{c}||}\cdot-\mathbf{n}_{i}$ (9)
@@
ranking views higher where the vertex normal $\mathbf{n}_{i}\in\mathbb{R}^{3}$ is parallel to camera ray defined by the cameras origin $\mathbf{o}_{c}$. The final camera view
$c^{\mathrm{opt}}_{f,i}=\arg\max_{c}(s_{f,c,i}*v_{f,c,i}),$ (10)
@@
is the maximum of the score times the vertex’ visibility $v_{f,c,i}\in\{0,1\}$ defining the final 3D correspondence point as $\tilde{\mathbf{P}}_{f,i}=\mathbf{P}_{f,c^{\mathrm{opt}}_{f,i},i}$.
Vertex Alignment Loss. Our template vertex correspondence loss
$\mathcal{L}_{\mathrm{corr-vertex}}(\mathbf{V}_{f})=\sum_{f=1}^{F}\sum_{i\in\mathcal{V}_{f}}||\mathbf{V}_{f,i}-\tilde{\mathbf{P}}_{f,i}||_{2}^{2}$ (11)
@@
only considers the set $\mathcal{V}_{f}$ of valid correspondences, i.e., pairs $(\mathbf{V}_{f,i},\tilde{\mathbf{P}}_{f,i})$ whose distance is smaller than $3cm$.
@@
Refined Training of the Drivable Template. Our drift-aware training objective is now a combination of Eq. 4 and Eq. 11, i.e., $\mathcal{L}_{\mathrm{template}}+\mathcal{L}_{\mathrm{corr-vertex}}$, which is used to train our drivable and depth-aligned template $M(\bm{\bar{\theta}}_{f},\mathbf{z}_{f})$. In practice, after optimizing the template mesh $\mathbf{V}_{f}$ via vertex-level alignment, we update the static texture $\mathbf{T}_{\mathrm{0}}$, recompute 2D correspondences using the (already) refined template, and perform a second round of vertex-level alignment. This cascaded refinement improves the coherency and accuracy of the template tracking, as evidenced by the qualitative and quantitative results in Sec. 4.2.
3.5 Texel Level Alignment

### v2

@@ 3.1. Gaussian-based Avatar Representation
Motion-aware Animatable Gaussian Textures. Building upon the drivable template mesh, we model the fine-grained and motion-aware appearance and geometry of the clothed human using Gaussian textures $\mathbf{T}^{\mathrm{gs}}_{f}\in\mathbb{R}^{N\times 62}$  (Pang et al., 2024) in the template’s UV space. Each of the $N$ texels covered by a triangle stores the parameters of a 3D Gaussian splat $(\boldsymbol{\bar{\mu}}^{\mathrm{uv}}_{i},\mathbf{\bar{d}}^{\mathrm{uv}}_{i},\mathbf{q}^{\mathrm{uv}}_{i},\mathbf{s}^{\mathrm{uv}}_{i},\mathbf{\alpha}^{\mathrm{uv}}_{i},\boldsymbol{\eta}^{\mathrm{uv}}_{i})_{f}\in\mathbb{R}^{62}$ where $i$ denotes the $i$th texel. Notably, the canonical Gaussian position $\boldsymbol{\bar{\mu}}_{\mathrm{uv},i}$ is derived from the non-rigidly deformed template mesh $\bar{\mathbf{V}}_{f}$ through barycentric interpolation. To model the finer-level details, i.e., wrinkles, of the dynamic clothed human, a learnable motion-aware offset $\mathbf{\bar{d}}_{\mathrm{uv},i}$ is applied to each Gaussian Splat in the canonical space. Similar to the drivable template mesh, the Gaussian splats can be posed from the deformed canonical position $(\boldsymbol{\bar{\mu}}_{i}+\mathbf{\bar{d}}_{\mathrm{uv},i})$ to the position $\boldsymbol{\mu}_{i}$ in posed space through Dual Quaternion skinning  (Kavan et al., 2007). The remaining parameters denote the rotation quaternion $\mathbf{q}^{\mathrm{uv}}_{i}$, anisotropic scaling $\mathbf{s}^{\mathrm{uv}}_{i}$, opacity $\mathbf{\alpha}^{\mathrm{uv}}_{i}$, and spherical harmonics coefficients $\boldsymbol{\eta}^{\mathrm{uv}}_{i}$.
@@
Predicting Motion-aware Gaussians. Thanks to the texel-based parameterization, the mapping between the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ and motion-aware Gaussian splats $\mathbf{T}^{\mathrm{gs}}_{f}\in\mathbb{R}^{N\times 62}$ can be formulated as an image-to-image translation task (Zhu et al., 2017). Specifically, the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ is represented using the positional $\mathbf{T}^{\mathrm{p}}_{f}$ and normal textures $\mathbf{T}^{\mathrm{n}}_{f}$ of the posed and deformed template $\mathbf{V}_{f}$. Then, two separate convolutional decoders (Pang et al., 2024) $\mathcal{E}_{\mathrm{geo}}(\mathbf{T}^{\mathrm{n}}_{f},\mathbf{T}^{\mathrm{p}}_{f})$ and $\mathcal{E}_{\mathrm{app}}(\mathbf{T}^{\mathrm{n}}_{f},\mathbf{T}^{\mathrm{p}}_{f})$ regress the geometry $(\mathbf{\bar{d}}^{\mathrm{uv}}_{i},\mathbf{q}^{\mathrm{uv}}_{i},\mathbf{s}^{\mathrm{uv}}_{i},\mathbf{\alpha}^{\mathrm{uv}}_{i})_{f}$ and appearance attributes $(\boldsymbol{\eta}^{\mathrm{uv}}_{i})_{f}$ of the 3D Gaussian Splats, respectively.
Splatting Gaussians to Image Space. We adopt Analytical Splatting (Liang et al., 2024)
(9) $I^{\mathrm{R}}_{f,c}=R^{\mathrm{ana}}(\mathbf{T}^{\mathrm{gs}}_{f},\mathbf{C}_{c})$
@@
to render the image $I^{\mathrm{R}}_{f,c}$ by integration over the actual pixel size during the splatting. This provides improved anti-aliasing over the original 3DGS tile-based rasterizer, which is particularly useful when supervising on ultra-resolution data, i.e., 6K resolution, and when focusing on detail preservation – the goal of this work.
@@
(10) $\mathcal{L}_{\mathrm{gau}}(I^{\mathrm{R}}_{f,c})=\mathcal{L}_{1}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c})+\mathcal{L}_{\mathrm{ssim}}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c}),+\mathcal{L}_{\mathrm{mrf}}(I^{\mathrm{R}}_{f,c},\mathbf{I}_{f,c})$
consisting of image-based $\mathcal{L}_{1}$, structural $\mathcal{L}_{\mathrm{ssim}}$, and perceptual losses $\mathcal{L}_{\mathrm{mrf}}$  (Wang et al., 2018).
3.2. Key Technical Challenge
@@
Misalignment in Depth. In the context of animatable clothed human avatars, the highly diverse clothing dynamics cannot be fully explained by skeletal motions alone, as similar skeletal poses may result in significantly different surface deformations, often referred to as the one-to-many mapping issue (Liu et al., 2021). As a result, when using Eq. 7 to supervise the avatar’s geometry (Eq. 1), the regression-based approach predicts an average surface that fits the multiple possible surface deformations in a least-squares sense due to this one-to-many mapping despite being supervised on accurate ground truth geometry. This manifests in misalignment in depth, which hinders accurate supervision of fine-grained appearance details, i.e., the motion-aware Gaussian Splats $\mathbf{T}^{\mathrm{gs}}_{f}$, since Gaussians may be wrongly projected onto training images, and therefore leads to blur or ghosting artifacts due to the conflicting supervision from different views (see Fig. 3 a).
On-surface Drift. Even if depth misalignment is resolved, the ground-truth geometry obtained from implicit-based methods (Wang et al., 2023) still lacks temporal correspondence among frames and correspondence between the template surface and the ground truth surface. Thus, the second source of error is noticeable drift on the surface between the drivable template mesh and the ground-truth surface since the Chamfer loss can be minimized even if the correspondence is wrong. For example in Fig. 3 b), the blue Gaussian is never mapped onto the correct location (marked in blue). Nonetheless, the Chamfer loss is minimized. Even worse, the Gaussian is mapped onto different ground truth surface locations for different frames. Thus, the rendering losses will assign different colors for different frames to the same Gaussian leading to a blurred average.
@@ 3.3. Depth Alignment
Figure 3. Key Technical Challenges. (a) The depth misalignment leads to gradient conflicts when supervised with multi-view images, like the blue splat. (b) Even if the depth misalignment is resolved, the surface drift between different frames results in gradient conflicts, resulting in averaged and blurry appearance.
@@
To address the misalignment in depth caused by the one-to-many mapping between the skeletal motion and the surface geometry, we adopt a learnable per-frame latent
(11) $\mathcal{H}_{\mathrm{lat}}(e(f/f_{\mathrm{max}}))=\mathbf{z}_{f}\in\mathbb{R}^{16}$
@@
as an additional input alongside the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ for predicting the drivable template mesh $\bar{\mathbf{V}}_{f}$. This latent accounts for the stochasticity and effectively establishes a one-to-one mapping. Here, $f_{\mathrm{max}}$ denotes the maximum number of training frames, $e(\cdot)$ is a positional encoding (Mildenhall et al., 2020), and $\mathcal{H}_{\mathrm{lat}}$ is a shallow MLP. We then update our deformable mesh (Eq. 1) by concatenating the learnable per-frame latent code $\boldsymbol{z}_{t}$ channel-wise to the input graph nodes of, both, the embedded deformation network $\mathcal{F}_{\mathrm{eg}}(\bar{\boldsymbol{\theta}}_{f},\mathbf{z}_{f})$ and the per-vertex deformation network $\mathcal{F}_{\mathrm{delta}}(\bar{\boldsymbol{\theta}}_{f},\mathbf{z}_{f})$.
@@
We refer to the drivable template mesh as $M(\boldsymbol{\bar{\theta}}_{f},\mathbf{z}_{f})$ – now also taking the learnable latent as input – and then train the drivable human avatar as before using the losses defined in Eq. 7. At test time, we set $\mathbf{z}_{f}$ to 0 to animate the character with novel motions, as it is intractable to predict the stochastic variations for unseen motions, and their influence on appearance and geometry is relatively minor compared to skeletal motion.
Moreover, to improve generalization to novel poses at the test time, inspired by  Li et al. (2024), we build a subspace $f_{\mathrm{pca}}$ of canonically deformed template meshes $\bar{\mathbf{V}}_{f}$ via Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak, 1993). To this end, the pose deformed template mesh under novel poses $\mathbf{V}^{\star}$ is computed through
(12) $\mathbf{V}^{\star}$ $=f_{\mathrm{dq}}(f_{\mathrm{pca}}(\bar{\mathbf{V}}),\boldsymbol{\theta}_{f})$
(13) $=f_{\mathrm{dq}}(\bar{\mathbf{V}}^{\star},\boldsymbol{\theta}_{f})$
where $\bar{\mathbf{V}}^{\star}$ is the canonical deformed template mesh in novel poses.
As an design alternative to compensate for stochastic effects of the clothed human that cannot be modeled by skeleton motions, Xiang et al. (2021) introduce latent channels for both coarse and fine-level geometry and appearance, while also minimizing the mutual information between the posed template and the latent channels. In contrast, we apply latent conditioning only to the coarse-level geometry, i.e., the template meshes $\bar{\mathbf{V}}_{f}$, whereas the finer-level geometry and appearance are conditioned solely on the coarse geometry without direct latent conditioning. This decomposition effectively models the stochasticity of large-scale clothing dynamics through the latent-conditioned template mesh, enabling the preservation of fine geometric and appearance details in subsequent stages of animatable Gaussian texture learning, while mitigating artifacts and jittering for novel poses.
3.4. Vertex-level Alignment
@@
Through the depth alignment (Sec. 3.3), the gap between the posed template mesh $\mathbf{V}_{f}$ and the ground truth surface $\mathbf{V}^{\mathrm{gt}}_{f}$ is significantly reduced by addressing the motion ambiguities. However, since the ground truth reconstruction $\mathbf{V}^{\mathrm{gt}}_{f}$ inherently lacks correspondence over time, surface drift is still inevitable, which will result in blurred appearance and geometry, as discussed in Sec. 3.2. Some prior works (Habermann et al., 2021a) propose to additionally supervise the geometric deformation with a differentiable image rendering loss to account for the drifting. However, we found that such loss is easily stuck in local minima while not being robust to shading and change in lighting (see also Sec. 4.3). Instead, our solution builds up on recent advances on foundational 2D point tracking methods (Karaev et al., 2024) that offer a promising alternative.
@@
2D Point Tracking. A 2D point tracker (Karaev et al., 2024) estimates 2D point tracks between a pair of images $(\mathbf{I},\mathbf{I}^{\prime})$, e.g., consecutive frames of a video. More formally, it can be defined as the function
(14) $h(\mathbf{I},\mathbf{I}^{\prime},\mathbf{p})=\mathbf{p}^{\prime}\in\mathbb{R}^{2}$
@@
taking the image pair as well as a 2D image coordinate $\mathbf{p}\in\mathbb{R}^{2}$ (of image $I$) and estimates the corresponding matching point $\mathbf{p}^{\prime}$ in image $I^{\prime}$. Since the point trackers are trained on large-scale real world datasets, they are typically robust to lighting changes and less likely to be trapped in local optima. However, our goal is to carefully align our drivable template $M(\boldsymbol{\bar{\theta}}_{f},\mathbf{z}_{f})$ over a long multi-view sequence, which is not natively supported by the original video point trackers.
@@
Render-to-image 2D Point Tracks. Our key idea is to compute 2D correspondence between the rendering $\mathbf{I}^{\mathrm{rast}}_{f,c}$ of the deformed template $\mathbf{V}_{f}$ observed from camera $c$ using a rasterizer (Laine et al., 2020) $R^{\mathrm{rast}}(\mathbf{V}_{f},\mathbf{T}_{\mathrm{0}},\mathbf{C}_{c})=\mathbf{I}^{\mathrm{rast}}_{f,c}$ and the ground truth training frame $\mathbf{I}_{f,c}$. Notably, as the deformed template suffers from surface drift, the deformable mesh rendering and the ground truth do not perfectly align and the point tracks $h(\mathbf{I}^{\mathrm{rast}}_{f,c},I_{f,c},\mathbf{p})$ $\mathbf{p}_{f,c,i}\in\mathbb{R}^{2}$ will capture exactly this drift. To acquire the static texture $\mathbf{T}_{\mathrm{0}}$ of the deformable mesh, we perform texture unprojection (Shetty et al., 2024) using the first frame of the multi-view video $\mathbf{I}_{0,c}$. Then, we can compute the drift for each vertex $\mathbf{V}_{f,i}\in\mathbb{R}^{3}$, i.e., the 2D point tracks $\mathbf{p}_{f,c,i}$, in image space as
(15) $\small\mathbf{p}_{f,c,i}$ $=h(\mathbf{I}^{\mathrm{rast}}_{f,c},\mathbf{I}_{f,c},\mathbf{C}_{c}\mathbf{V}_{f,i})$
(16) $=h(\mathbf{I}^{\mathrm{rast}}_{f,c},\mathbf{I}_{f,c},\mathbf{v}_{f,c,i})$
by querying $h(\cdot,\cdot,\cdot)$ at the vertices’ projected 2D position $\mathbf{v}_{c,f,i}\in\mathbb{R}^{2}$.
@@
3D Lifting and Consensus Finding. Next, we lift per-view 2D correspondences $\{\mathbf{p}_{f,c,i}\}_{c=1}^{C}$ of vertex $i$ to 3D correspondences $\{\mathbf{P}_{f,c,i}\}_{c=1}^{C}$ by querying depth maps followed by applying the inverse camera transformation:
(17) $\mathbf{P}_{f,c,i}=\mathbf{C}^{-1}_{c}\mathbf{N}_{f,c}[\mathbf{p}_{f,c,i}],$
where $\cdot$] denotes bilinear interpolation. To find consensus across views, we define a score function
[…]
@@
ranking views higher where the vertex normal $\mathbf{n}_{i}\in\mathbb{R}^{3}$ is parallel to camera ray defined by the cameras origin $\mathbf{o}_{c}$. The optimal camera view
(19) $c^{\mathrm{opt}}_{f,i}=\arg\max_{c}(s_{f,c,i}*v_{f,c,i}),$
@@
is the maximum of the score times the vertex’ visibility $v_{f,c,i}\in\{0,1\}$ defining the final 3D correspondence point as
(20) $\tilde{\mathbf{P}}_{f,i}=\mathbf{P}_{f,c^{\mathrm{opt}}_{f,i},i}.$
Vertex Alignment Loss. Our template vertex alignment loss
[…]
@@
sums over all frames while only considering the set $\mathcal{V}_{f}$ of valid correspondences, i.e., pairs $(\mathbf{V}_{f,i},\tilde{\mathbf{P}}_{f,i})$ whose distance is less than $3cm$.
Refined Training of the Drivable Template. Our drift-aware training objective
(22) $\mathcal{L}_{\mathrm{template}}+\mathcal{L}_{\mathrm{cor-vrt}}$
@@
is now a combination of Eq. 7 and Eq. 21, which is used to train our drivable and depth-aligned template $M(\boldsymbol{\bar{\theta}}_{f},\mathbf{z}_{f})$. In practice, after optimizing the template mesh $\mathbf{V}_{f}$ via vertex-level alignment, we update the static texture $\mathbf{T}_{\mathrm{0}}$, recompute 2D correspondences using the (already) refined template, and perform a second round of vertex-level alignment. This cascaded refinement improves the coherency and accuracy of the template tracking, as evidenced by the qualitative and quantitative results in Sec. 4.3.
3.5. Texel-level Alignment